# Sample Observed Effects: Enumeration, Randomization and Generalization

### v5

@@ 5.1 Randomization and Parametric Inference
We will use the notation often employed in Structurual Causal Models (SCM) to relate the previous concepts to research in parametric effect inference. In SCM models, variables are typically specified by an unknown function of a set of endogeneous variables and an exogeneous error. We have, similarly, specified DGPs $Z=\{X,U\}$ with a collection of exogenous factors $U$, in addition to the observed $X$. Factors in $U$ are unobserved and not affected by $X$, but can affect $X$. More specifically, a SCM specifies a collection of functions $f_{a}$ indexed by endogenous variables, $f=(f_{a}:a,b,c,...)$. The SCM also calls for the specification of the relation between the random variables $U$ and $X$. This is done through a set of functions $f=(f_{W},f_{a},f_{y})$, where $A$ is a random indicator for a treatment $a$, $W$ are the other (non-treated) in-sample factors, $W=X-\{a\}$, and $y$ is the outcome of interest. The values of $W$, $A$, and $Y$ are then deterministically assigned by $U$, according to  [van der Laan and Rose, 2011]
@@
Randomized controlled trials (RCTs) are one type of controlled experiment where subjects are randomized to receive a specific treatment. For example, when each subject is assigned to a treatment based on the flip of a coin. In this case, the differences between the outcomes of treatment and control groups become are due to treatment, as all other factors would be balanced (free of selection bias), up to random error. In an RCT, the treatment assignment mechanism, $P(A=1|W)$, is known (e.g., constant and $0.5$ as in the coin flip). This is different from non-experimental data where treatment, or treatment exposure, is not assigned. A RCT can be, nonetheless, written as in Eq.(16-17,19) in SCM notation,
@@
The only difference of Eq.(15) from Eq.(16-17,19) is the structural equation for $f_{a}$ in Eq.(17). The baseline variables $W$ play no role in the generation of $A$. These assumptions can also be illustrated with causal graphs, Fig.2(c). Since $a$ is randomized, this implies that $U_{a}$ is independent of $U_{y}$ and $U_{W}$, and, in this case, no arrows connect $U_{a}$ to $U_{y}$ and $U_{a}$ to $U_{w}$ in the graphical model specification. Eq.(LABEL:eq-rctthis) reviews the conditions on sample gap distributions proposed in this article and in Eq.(8).
@@ 5.2 Non-parametric Effect Estimation
$\hat{\Delta y(a)}_{n}=\arg\min_{\theta}\left\{\frac{1}{m-1}\sum_{w\in X-\{a\}}L\left(\Delta y(a),Q,P_{w}(a)\right)\right\}$
Here, $L$ is the loss function over hyperparameters $\theta$ that quantifies the discrepancy between the true parameter $\Delta y(a)$, the data-generating distribution $Q$, and the empirical distribution $P_{w}$ induced by a specific treatment assignment. The optimization is typically performed using ensemble or black-box machine learning techniques. The optimization thus aims to find hyperparameters $\theta$ that minimizes the loss across all treatment assignment mechanisms, resulting in an estimator that is robust to different ways treatments may have been assigned in the non-experimental sample. The approach here has the same goal, but relies on a theoretical connection between treatment-background assignments, in one hand, and their relation to generalizability and confoudness of effects on the other. The theory provides an answer to when specific effect observations can be said to have the later properties, conditional on a fixed mechanism (as identified by sample gap statistics and tests). We seek to demonstrate that under specific mechanisms samples contain effects that are externally valid. The approach holds a similar relationship to parametric methods that employ black-box techniques across time to generate treatment-control pairs for effect estimation [Calonico et al., 2019, Abadie, 2021].
In respect to outcome functions, Eq.(19), a general non-parametric strategy is to condition outcomes on all in-sample factors and take their difference in means in each stratum. This leads to an estimate of the effect of treatment as the average over all strata. This approach is significantly more efficient than an unadjusted treatment effect estimate. That’s because the Law of Total Probabilities allows such methods to control (in-sample) confounding without the assumption that effects are constant across strata. If $q=0$, the g-formula  [Keil et al., 2014] implements this idea. For example, an outcome could be expressed as
@@
which indicates that we are summing over each possible value of $W=X-\{a\}$, and $P(W=w)$ is the probability that $W$ takes on the value $w$ in the reference population when only $a$ is left to vary. This adjusted estimator of the effect of $a$ is a nonparametric maximum likelihood estimator for $Y$ for a single treatment  [Keil et al., 2014, van der Laan and Rose, 2011]. Drawbacks are the requirement that $q=0$ and the loss in performance in strata with small number of sample units. Analogously, in Eq.(3) effects are conditioned on all their backgrounds (where order is also taken to matter). The strategy in Eq.(21) is often combined with conditional independence tests to identify plausible in-sample confounders. While we presently favor simplicity to highlight theoretical considerations and unobserved confounding, elements of these approaches (such as further in-sample conditional independence tests of the g-formula and the supervised plug-in flexibility of targeted learning) could be further combined with the proposed.
@@ 7 Experiments
These correspond to out-of-sample prediction of outcomes, prediction of intervention effects, and marginal factor effect estimation (evaluated against the ground-truth in simulations). The first is the central problem in supervised Machine Learning, the second is related to effect estimation with counterfactual and matching estimators [Morgan and Winship, 2007], and the third to both causal effect estimation [Pearl, 2000], and supervised explainability  [Burkart and Huber, 2021]. Effect estimation is, in the present perspective, an extreme omitted-variable problem. The problem becomes easier with decreases in the number $q$ of unobserved factors and the proposed framework can also be used to study the performance of generic samples and models.
@@
The last number indicates that, to form a partial permutation, we select $d$ unique factors to fix and be organized in $\binom{m}{d}$ ways, each with $D_{m-d}$ possible disjoint orderings of the non-selected factors. Fig.3(a) shows the number of observed permutations present on increasingly larger samples (balanced). Each (green) curve corresponds to the number of permutations in samples of size $n=\{200,350,500,1000\}$ (blue curves show the number of squares). The lower panel shows combinatorial limits (red) for the number of partial permutations and squares according to Eq.(23), $m=10$. The figure illustrates, in particular, the relationship between sample sizes $n$ and the subset of permutations $\Pi_{n}(Z)$ that effects for factors $Z$ can be asymptotically observed under222since $q\leq 20$ in these illustrations all permutations in a sample can be enumerated with the approach in Appendix.I, for each number $d$ of fixation-points, $d\leq m$. The relationship between the DGP and enumerable permutations or their speed of enumeration was, for example, formulated in Eq.(2) and illustrated in Fig.1(b). We expect these quantities to impact sample biases and predictive performance, Eq.(3), irrespective of the DGP. The same is shown as histograms of square member counts, Fig.3(b). At $n{=}200$, no differences of size larger than $4$ are observed. A full square of size $10$ is completely observed only at $n{=}1024$ (bottom-right) in the balanced case, and $n\approx 20K$ in the unbalanced case. This was anticipated by Eq.(22) and follow from simple sample ordinal statistics. The first sample size is $\tilde{n}=2^{9+1}=1024$. The latter sample has rare factors, ${\rm I\kern-1.99997ptE}[min_{a\in X}(p(a))]=0.025$, which leads to an approximate sample requirement of $\tilde{n}=2^{9}/0.025=20480$. Similar sample size recommendations will be indicated for each of the increasingly complex simulated and real-world cases considered below.
@@ 7.1 Outcome Prediction
An understanding of the relationship among samples and estimators’ ACC and EV is necessary to devise accurate and maximally general subsamples. To that end, we consider ACC (percentage of correctly classified cases in sample validation sections) as we increase $n$. To that end, we divide samples in two random sections: internal and external. The internal section is divided in typical training and validation subsections. We report accuracy of cross validation in the internal section (green) and external (red), with $4$ folds across cases. When ordered arbitrarily, or randomly, the internal-external division is inconsequential (as long as individual sections contain enough samples to cross algorithm-specific overfitting thresholds). We will define, however, alternative sample transversal orders, and consider how they impact learning performance. Starting with $n{=}1$, we train several algorithms with increasing $n$, and under different orders. With each unitary increase of $n$, an observation in the external section is transferred to the internal, and the problem of model and effect generalization becomes easier. Accordingly, a sample that can generalize its model and effects to external populations the earliest can be said to have higher EV. We use: three pre-specified XGBoost GBM (Gradient Boosting Machine) models, a grid of GLMs (Generalized Linear Model), a Random Forest (DRF), LASSO and Ridge regressions, five pre-specified GBMs, a near-default Deep Neural Net, an Extremely Randomized Forest (XRT), a random grid of XGBoost GBMs, a random grid of GBMs, and a random grid of Deep Neural Nets, as well as Stacked Ensembles with all previous models. Searched and optimized parameters in each case are listed in Appendix.F.
@@ 7.3 Causal Effect Estimation and Black-box Explanations
The approach here also establishes a connection to variance-based sensitivity analysis. Sensitivity analysis decompose systems’ error to estimate individual factor contributions. The Sobol index  [Sobol, 2001] is defined from the conditional variance of a factor, $\nicefrac{{\text{VAR}{\rm I\kern-1.39998ptE}[y|x_{i}]}}{{\text{VAR}(y)}}=\nicefrac{{\text{VAR}(\beta)}}{{\text{VAR}(y)}}$, where $\beta$ are effect estimates (often from Analysis of Variance, ANOVA). Most explaining and sensitivity analysis approaches start with a functional for outcomes, $y=f(a,b,c,...)$ where input factors $a,b,c,...$ are i.i.d. Black-box explainers subject boxes (algorithms) to a large number of variations, while sensitivity analysis decomposes their errors. The goal for both is to indicate the importance of individual input variables, especially in respect to sample predictive performance. The Shapley value comes from coalitional game theory and was devised axiomatically. Beyond its strategic origins, Shapley’s axioms are important requirements to decompose changes in any system  [de Boer and Rodrigues, 2020]. As it is often used with random sampling, it is closely related to bootstrap and permutation methods for black-box regressions.
Squares were used here as means to represent sets of effect observations for a unit $x_{0}$, their permutations in samples, and, finally, to define non-parametric statistics for sample factors’ effects. Our initial definition of effects, Eq.(3), led to consequent definitions and interpretations for the EV, Definition (2), and CF, Definition (3), of counterfactual observations and samples. The unit of analysis here is $\Delta y$, an outcome difference observation, and not the outcomes of regressions. Permutations are not random sample intake orders, but ’observed’ permutations in samples that fulfilled a set of combinatorial and gap conditions, and offer additional guarantees over effect estimates. Eq.(3) thus uses the complete set of $m$-way observed differences in a sample to permute observed counterfactual outcomes and derive effect estimates. The full set of effect background variation, ${\Pi}_{n}(Z\/{a/})$, is the full set of EV-increasing sequences in samples (or their ’external variation’) for individual sample populations.
@@ 7.5 The 80/20 Sample Split
The most universally used rule-of-thumb in ML practice is the 80/20 ratio when splitting samples into train and test subsamples. Although it has remained at the theoretical margins of ML research, largely as application to the problem of overtraining in neural networks, a natural question is: why this proportion? There are two common answers. The first is heuristic, as a consequence of the Pareto Principle. The principle states that ’80% of effects come from 20% of causes’  [Chen et al., 1994]. Causes are used here vaguely, but the principle explains many natural and artificial phenomena. The second answer is more precise, formulated as a scaling-law for the ratio  [Guyon, 1997, Bahri et al., 2021], generalizing  [Amari et al., 1995, Kearns, 1995]. They find that ’the fraction of patterns reserved for the validation set should be inversely proportional to the square root of the number of free adjustable parameters’. In essence, the optimal split is therefore determined by the number of unique factors in a sample, and not its gross number of observations. Eq.(22)(multiple-squares) derived simple bounds, $\tilde{n}$, for sample sizes. The pareto principle and ratio scaling-law offer similar recommendations. Fig.7(o) shows mean difference between recommended sizes, $\tilde{n}$, and the optimal split, $n_{opt}$, across all previous simulation cases. In each case, a sample with $n^{\prime}>100$ observations is divided into 100 random and increasing subsamples. Each subsample adds $n^{\prime}/100$ new observations to the previous and is used as training for the previous supervised methods. The optimal split $n_{opt}$ is taken as the subsample size with an inflection point in the best performance (ACC) across all methods (all samples had ACC peaks, if non-unique the smallest size was taken as optimal). Error is thus $(\tilde{n}-n_{opt})/n^{\prime}$. Although an asymptotic approximation, relating to the asymptotic number of permutations in unbalanced samples, Eq.(22) describes well trade-offs in relation to the EV of training samples.
@@ 8 Discussion and Conclusion
The central methodological challenge in the Sciences, Policy-making and Design remains the evaluation of counterfactual statements (Did this treatment caused the result of interest? Did this policy?) The counterfactual definition of effects  [Rubin, 2005, Morgan and Winship, 2007] formulate sample properties necessary for effects to be free of selection biases, Eq.(1). It is based on the fixation of sample effect backgrounds. The strategy in this work has been, in contrast, to focus on conditions where the amount of unobserved variation is guaranteed to be large (sampling is EV-increasing) and shared among factors (CF non-increasing). This alternative definition, as effects that remain invariant in the face of large external variation, can lead, instead, to generalizeable effect estimates. Under these conditions, causes can also be separated from confounders using effect variance estimates, Principle (1). These considerations led to the notion of effect background randomization, and sample conditions under which effects become sufficiently randomized. The perspective also allowed us to reconsider problems that have proven difficult for the standard counterfactual formulation, such as their sample size requirements and unobserved variable biases.
@@
An advantage of the proposed approach is thus bringing causal and predictive concepts closer. We demonstrated generalizability bounds for popular supervised predictors and black-box explainers [Burkart and Huber, 2021]. Supervised prediction, black-box explanation and effect estimation approaches should state, side-by-side with their estimates, the conditions on which they are expected to hold (sample sizes, correlations, completeness, etc.). This is unknown for most current causal effect estimation solutions, IME explainers, and most out-of-sample prediction solutions. Until these are stated, it is difficult to, in practice, trust their outputs. The approach allows us to model effects in both linear and non-linear data without the, rarely know and difficult to infer, parametric models for outcomes or strict independence among factors. We discussed out-of-sample, counterfactual, correlated and omitted variables prediction, as well as causal effect estimation, in simulations and an important real-world example. Correlated and counterfactual prediction observed, in particular, significant accuracy gains. A central result demonstrated here is that the EV of interventions are, to some extent, predictable from combinatorial properties of the populations they act upon. To a broad audience, Machine Learning methods are inherently limited due to their lack of generalizability and selection bias guarantees, when compared to, for example, experimental hypothesis testing. We believe the work could help build further connections between predictive and causal techniques.

### v6

@@ 5.1 Randomization and Parametric Inference
We will use the notation often employed in Structural Causal Models (SCM) to relate the previous concepts to research in parametric effect inference. In SCM models, variables are typically specified by an unknown function of a set of endogeneous variables and an exogeneous error. We have, similarly, specified DGPs $Z=\{X,U\}$ with a collection of exogenous factors $U$, in addition to the observed $X$. Factors in $U$ are unobserved and not affected by $X$, but can affect $X$. More specifically, a SCM specifies a collection of functions $f_{a}$ indexed by endogenous variables, $f=(f_{a}:a,b,c,...)$. The SCM also calls for the specification of the relation between the random variables $U$ and $X$. This is done through a set of functions $f=(f_{W},f_{a},f_{y})$, $A$ is a random indicator for a treatment $a$, $W$ are the other (non-treated) in-sample factors, $W=X-\{a\}$, and $y$ is the outcome of interest. The values of $W$, $A$, and $Y$ are then deterministically assigned by $U$, according to  [van der Laan and Rose, 2011]
@@
Randomized controlled trials (RCTs) are one type of controlled experiment where subjects are randomized to receive a specific treatment. For example, when each subject is assigned to a treatment based on the flip of a coin. In this case, differences between the outcomes of treatment and control groups indicate the treatment’s effect, as all other factors would be balanced (free of selection bias), up to a random error. In an RCT, the treatment assignment mechanism, $P(A=1|W)$, is known (e.g., constant and $0.5$ in the coin flip). This is different from non-experimental data where treatment, or treatment exposure, is not assigned. A RCT can be, nonetheless, written in SCM notation as in Eq.(15-16,18),
@@
The only difference between Eq.(14) and Eq.(15-16,18) is the structural equation for $f_{a}$ in Eq.(16). The baseline variables $W$ play no role in the generation of $A$. These assumptions can also be illustrated with causal graphs, Fig.2(c). Since $a$ is randomized, this implies that $U_{a}$ is independent of $U_{y}$ and $U_{W}$, and, in this case, no arrows connect $U_{a}$ to $U_{y}$ and $U_{a}$ to $U_{w}$ in the graphical model. Eq.(LABEL:eq-rctthis) reviews the conditions on sample gap distributions proposed in this article and in Eq.(8).
@@ 5.2 Non-parametric Effect Estimation
$\hat{\Delta y(a)}_{n}=\arg\min_{\theta}\left\{\frac{1}{m-1}\sum_{w\in X-\{a\}}L_{\theta}\left(\Delta y(a),Q,P_{w}(a)\right)\right\}$
Here, $L_{\theta}$ is the loss function over hyperparameters $\theta$ that quantifies the discrepancy between the true parameter $\Delta y(a)$, the data-generating distribution $Q$, and the empirical distribution $P_{w}$ induced by a specific treatment assignment. The optimization is typically performed using ensemble or black-box machine learning techniques. The optimization thus aims to find hyperparameters $\theta$ that minimizes the loss across all treatment assignment mechanisms, resulting in an estimator that is robust to different ways treatments may have been assigned in the non-experimental sample. The approach here has the same goal, but relies on a theoretical connection between treatment-background assignments, in one hand, and their relation to generalizability and confoundness of effects on the other. The theory provides an answer to when specific effect observations can be said to have the later properties, conditional on a fixed mechanism (as identified by sample gap statistics and tests). We seek to demonstrate that, under specific mechanisms, samples contain effects that are externally valid. The approach holds a similar relationship to parametric methods that employ black-box techniques across time to generate treatment-control pairs for effect estimation [Calonico et al., 2019, Abadie, 2021].
In respect to outcome functions, Eq.(18), a general non-parametric strategy is to condition outcomes on all in-sample factors and take their difference in means in each resulting stratum. This leads to an estimate of the effect of treatment as the average over all strata which is significantly more efficient than an unadjusted treatment effect estimate. That’s because the Law of Total Probabilities allows such methods to control (in-sample) confounding without the assumption that effects are constant across strata. If $q=0$, the g-formula  [Keil et al., 2014] implements this idea. For example, an outcome could be expressed as
@@
which indicates that we are summing over each possible value of $W=X-\{a\}$, and $P(W=w)$ is the probability that $W$ takes on the value $w$ in the reference population when only $a$ is left to vary. This adjusted estimator of the effect of $a$ is a nonparametric maximum likelihood estimator for $Y$ for a single treatment  [Keil et al., 2014, van der Laan and Rose, 2011]. Drawbacks are the requirement that $q=0$ and the loss in performance in strata with a small number of units. Analogously, in Eq.(3) effects are conditioned on all their backgrounds (where order is also taken to matter). The strategy in Eq.(20) is often combined with conditional independence tests to identify plausible in-sample confounders. While we presently favor simplicity to highlight theoretical considerations and unobserved confounding, elements of both these approaches (such as further in-sample conditional independence tests of the g-formula and the supervised plug-in flexibility of targeted learning) could be further combined with the proposed.
@@ 7 Experiments
These correspond to out-of-sample prediction of outcomes, prediction of intervention effects, and marginal factor effect estimation (evaluated against the ground-truth in simulations). The first is the central problem in supervised Machine Learning, the second is related to effect estimation with counterfactual and matching estimators [Morgan and Winship, 2007], and the third to both causal effect estimation [Pearl, 2000] and supervised explainability  [Burkart and Huber, 2021]. Effect estimation is, in the present perspective, an extreme omitted-variable problem. The problem becomes easier with decreases in the number $q$ of unobserved factors and the proposed framework can thus also be used to study the performance of generic samples and models.
@@
The last number indicates that, to form a partial permutation, we select $d$ unique factors to fix and be organized in $\binom{m}{d}$ ways, each with $D_{m-d}$ possible disjoint orderings of the non-selected factors. Fig.3(a) shows the number of observed permutations present on increasingly larger samples (balanced). Each (green) curve corresponds to the number of permutations in samples of size $n=\{200,350,500,1000\}$ (blue curves show the number of squares). The lower panel shows combinatorial limits (red) for the number of partial permutations and squares according to Eq.(22), $m=10$. The figure illustrates, in particular, the relationship between sample sizes $n$ and the subset of permutations $\Pi_{n}(Z)$ that effects for factors $Z$ can be asymptotically observed under222since $q\leq 20$ in these illustrations all permutations in a sample can be enumerated with the approach in Appendix.I, for each number $d$ of fixation-points, $d\leq m$. The relationship between the DGP and enumerable permutations or their ’speed of enumeration’ was, for example, formulated in Eq.(2) and illustrated in Fig.1(b). We expect these quantities to impact sample biases and predictive performance, Eq.(3), irrespective of the DGP. The same is shown as histograms of square member counts, Fig.3(b). At $n{=}200$, no differences of size larger than $4$ are observed. A full square of size $10$ is completely observed only at $n{=}1024$ (bottom-right) in the balanced case, and $n\approx 20K$ in the unbalanced case. This was anticipated by Eq.(21) and follow from simple sample ordinal statistics. The first sample size is $\tilde{n}=2^{9+1}=1024$. The latter sample has rare factors, ${\rm I\kern-1.99997ptE}[min_{a\in X}(p(a))]=0.025$, which leads to an approximate sample requirement of $\tilde{n}=2^{9}/0.025=20480$. Similar sample size recommendations will be indicated for each of the increasingly complex simulated and real-world cases considered below.
@@ 7.1 Outcome Prediction
An understanding of the relationship among samples and estimators’ accuracy and EV is necessary to devise accurate and maximally general subsamples. To that end, we consider ACC (percentage of correctly classified cases in sample validation sections) as we increase $n$. We divide samples in two random sections: internal and external. The internal section is divided in typical training and validation subsections. We report accuracy of cross validation in the internal section (green) and external (red), with $4$ folds across cases. When ordered arbitrarily, or randomly, the internal-external division is inconsequential (as long as individual sections contain enough samples to cross algorithm-specific overfitting thresholds). We will define, however, alternative sample transversal orders, and consider how they impact learning performance. Starting with a small $n$, we train several algorithms with increasing $n$, and under different orders. With each unitary increase of $n$, an observation in the external section is transferred to the internal, and the problem of model and effect generalization becomes easier. Accordingly, a sample that can generalize its model and effects to external populations the earliest can be said to have higher EV. We use: three pre-specified XGBoost GBM (Gradient Boosting Machine) models, a grid of GLMs (Generalized Linear Model), a Random Forest (DRF), LASSO and Ridge regressions, five pre-specified GBMs, a near-default Deep Neural Net, an Extremely Randomized Forest (XRT), a random grid of XGBoost GBMs, a random grid of GBMs, and a random grid of Deep Neural Nets, as well as Stacked Ensembles with all previous models. Searched and optimized parameters in each case are listed in Appendix.F.
@@ 7.3 Causal Effect Estimation and Black-box Explanations
The approach here also establishes a connection to variance-based sensitivity analysis  [Bénesse et al., 2024]. Sensitivity analysis decompose systems’ error to estimate individual factor contributions. The Sobol index  [Sobol, 2001] is defined from the conditional variance of a factor, $\nicefrac{{\text{VAR}{\rm I\kern-1.39998ptE}[y|x_{i}]}}{{\text{VAR}(y)}}=\nicefrac{{\text{VAR}(\beta)}}{{\text{VAR}(y)}}$, where $\beta$ are effect estimates (often from Analysis of Variance, ANOVA). Most explaining and sensitivity analysis approaches start with a functional for outcomes, $y=f(a,b,c,...)$ where input factors $a,b,c,...$ are i.i.d. Black-box explainers subject boxes (algorithms) to a large number of variations, while sensitivity analysis decomposes their errors. The goal for both is to indicate the importance of individual input variables, especially in respect to sample predictive performance. The Shapley value comes from coalitional game theory and was devised axiomatically. Beyond its strategic origins, Shapley’s axioms are important requirements to decompose changes in any system  [de Boer and Rodrigues, 2020]. As it is often used with random sampling, it is closely related to bootstrap and permutation methods for black-box regressions.
Squares were used here as means to represent sets of effect observations for a unit $x_{0}$, their permutations in samples, and, finally, to define non-parametric statistics for sample factors’ effects. Our initial definition of effects, Eq.(3), led to consequent definitions and interpretations for the EV, Definition (2), and CF, Definition (3), of counterfactual observations and samples. The unit of analysis here is $\Delta y$, an outcome difference observation, and not the outcomes of regressions. Permutations are not random sample intake orders, but ’observed’ permutations in samples that fulfilled a set of combinatorial and gap conditions, and offer additional guarantees over effect estimates. Eq.(3) thus uses the complete set of $m$-way observed differences in a sample to permute observed counterfactual outcomes and derive effect estimates. The full set of effect background variation, $\Pi_{n}(Z/\{a\})$, is the full set of EV-increasing sequences in samples (or their ’external variation’) for individual sample populations.
@@ 7.5 The 80/20 Sample Split
The most universally used rule-of-thumb in ML practice is the 80/20 ratio when splitting samples into train and test subsamples. Although it has remained at the theoretical margins of ML research, largely as application to the problem of overtraining in neural networks, a natural question is: why this proportion? There are two common answers. The first is heuristic, as a consequence of the Pareto Principle. The principle states that ’80% of effects come from 20% of causes’  [Chen et al., 1994]. Causes are used here vaguely, but the principle explains many natural and artificial phenomena. The second answer is more precise, formulated as a scaling-law  [Guyon, 1997, Bahri et al., 2021], generalizing  [Amari et al., 1995, Kearns, 1995]. They find that ’the fraction of patterns reserved for the validation set should be inversely proportional to the square root of the number of free adjustable parameters’. In essence, the optimal split is therefore determined by the number of unique factors in a sample, and not its gross number of observations. Eq.(21)(multiple-squares) derived simple bounds, $\tilde{n}$, for sample sizes. The pareto principle and ratio scaling-law offer similar recommendations. Fig.7(o) shows mean difference between recommended sizes, $\tilde{n}$, and the optimal split, $n_{opt}$, across all previous simulation cases. In each case, a sample with $n^{\prime}>100$ observations is divided into 100 random and increasing subsamples. Each subsample adds $n^{\prime}/100$ new observations to the previous and is used as training for the previous supervised methods. The optimal split $n_{opt}$ is taken as the subsample size with an inflection point in the best performance (ACC) across all methods (all samples had ACC peaks, if non-unique the smallest size was taken as optimal). Error is thus $(\tilde{n}-n_{opt})/n^{\prime}$. Although an asymptotic approximation, relating to the asymptotic number of permutations in unbalanced samples, Eq.(21) describes well trade-offs in relation to the EV of training samples.
@@ 8 Discussion and Conclusion
The central methodological challenge in the Sciences, Policy-making and Design remains the evaluation of counterfactual statements (Did this treatment caused the result of interest? Did this policy?) The counterfactual definition of effects  [Rubin, 2005, Morgan and Winship, 2007] formulate sample properties necessary for effects to be free of selection biases, Eq.(1). It is based on the fixation of effect backgrounds in samples. The strategy in this work has been, in contrast, to focus on conditions where the amount of unobserved variation is guaranteed to be large (sampling is EV-increasing) and shared among factors (CF non-increasing). This alternative definition, as effects that remain invariant in the face of large external variation, can lead, instead, to generalizeable effect estimates. Under these conditions, causes can also be separated from confounders using effect variance estimates, Principle (1). These considerations led to the notion of effect background randomization, and sample conditions under which effects become sufficiently randomized. The perspective also allowed us to reconsider problems that have proven difficult for the standard counterfactual formulation, such as their sample size requirements and unobserved variable biases.
@@
An advantage of the proposed approach is thus bringing causal and predictive concepts closer. We demonstrated generalizability bounds for popular supervised predictors and black-box explainers [Burkart and Huber, 2021]. Supervised prediction, black-box explanation and effect estimation approaches should state, side-by-side with their estimates, the conditions on which they are expected to hold (sample sizes, correlations, completeness, etc.). This is unknown for most current causal effect estimation solutions, IME explainers, and most out-of-sample prediction solutions. Until these are stated, it is difficult to, in practice, trust their outputs. The approach allows us to model effects in both linear and non-linear data without the, rarely know and difficult to infer, parametric models for outcomes or independence among factors. We discussed out-of-sample, counterfactual, correlated and omitted variables prediction, as well as causal effect estimation, in simulations and an important real-world example. Correlated and counterfactual prediction observed, in particular, significant accuracy gains. A central result demonstrated here is that the EV of interventions are, to some extent, predictable from combinatorial properties of the populations they act upon. To a broad audience, Machine Learning methods are inherently limited due to their lack of generalizability and selection bias guarantees, when compared to, for example, experimental hypothesis testing. We believe the work could help build further connections between predictive and causal techniques.